Please kind person, I need your help. Can you please assist me with this question?

Perform the operation 1/2, (mod 5)
Select correct answer
1/2 = 2
1/2 = 11
1/2 = 3

Answer by jim_thompson5910(35256)

1/2 is really 2^(-1), which is the multiplicative inverse of 2 (mod 5)

So we need to find a number that when you multiply it by 2, you'll get a result that is congruent to 1 (mod 5)


Notice how 2*3 = 6 = 1 (mod 5)

So 2^(-1) = 3

Using your notation of 1/2, this means 1/2 = 3 (mod 5)